f(x,y) = y^3 + ysin2x + e^(x+y) is diffrentiable at (1,-1).
1
Expert's answer
2020-06-15T19:35:10-0400

"f(x,y) = y^3 + ysin2x + e^{x+y}";

"f(x,y)" is differential at (a,b) if f is continuous at (a,b) and the partial derivatives "\frac{\partial f}{\partial x}" and "\frac{\partial f}{\partial y}" are both defined and continuous at (a,b).

f is continuous as a sum of continuous functions.

Let's find partial derivatives of f.

"\frac{\partial f}{\partial x}=2ycos2x+e^{x+y}."

"\frac{\partial f}{\partial y}=2y^2+sin2x+e^{x+y}."

The partial derivatives are both defined at (-1,1) and continuous as the sums of continuous functions.

Hence "f(x,y)=y^3+ysin2x+e^{x+y}" is differettiable at (-1; 1)